Circular No. 7851

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 C1 (IKEYA-ZHANG) M. Fujii, Fujii Bisei Observatory (FBO); K. Ayani, Bisei Astronomical

Observatory (BAO); and H. Kawakita, Gunma Astronomical Observatory, report that low-resolution spectra of comet C/2002 C1 obtained with the FBO 0.28-m reflector on Mar. 3.41 UT show the Na I emission line at 589.3 nm. Confirming spectroscopy at higher resolution (0.05 nm) with the BAO 1.01-m reflector on Mar. 6.41 resolved the Na I D_{1} and D_{2} emission lines. The stronger D_{2} line has an equivalent width of 0.5 nm and is blueshifted by −33 km/s, which is consistent with the geocentric velocity of the comet.

SUPERNOVA 2002bn IN UGC 6527 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002bn (cf. IAUC 7846), obtained by V. Hradecky on Mar. 9.28 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, four weeks after maximum.

SUPERNOVA 2002bo IN NGC 3190 R. Chornock, W. D. Li, and A. V. Filippenko, University of Califor-

nia, Berkeley, report that inspection of CCD spectra (range 330–1000 nm) obtained on Mar. 11 UT with the Shane 3-m reflector at Lick Observatory reveals that SN 2002bo is a supernova of type Ia. The Si II absorption near 585 nm is unusually strong, perhaps implying that the supernova is somewhat underluminous. However, the spectrum lacks the strong Ti II features typical of SN 1991bg-like objects (Filippenko et al. 1992, A.J. 104, 1543). The expansion speed of the ejecta, measured from the minimum in the Si II 635.5-nm trough, is ∼ 16 000 km/s; such a high value suggests that the object is quite young.

SUPERNOVA 2002bl IN UGC 5499 D. Rich, Hampden, ME, reports a prediscovery observation of SN

2002bl (cf. IAUC 7845) at mag ∼ 17.8 in an unfiltered CCD image taken with a 0.26-m telescope on Feb. 12.21 UT.

SUPERNOVA 2002bq IN MCG –06-21-10

On IAUC 7848, the MCG number was given incorrectly.

2002 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green